Shift report 20011106 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Nov 06 2001 - 07:47:08 EST

  • Next message: Ian Bearden: "FW: QM2001 Proceedings"

    ------------------------------------------------------------
    Date: Tue Nov  6 07:47:08 2001, Shift: 00:00-08:00
    Supervisor: BHS, with: 
    Log book#10 pages 100 - 105
    
    Shift Summary:
    ------------------------------------------------------------
    1: A good shift :-) We've had one big and stable store 
    lasting from 02.30 until 07.05, dumped by votes from
    BRAHMS and PHENIX.
    
    2: Runs
    Run    Trig6   Setting
    5681    124k   MRS@90, D5@2B, FS@12, D1-D4@1/8A
    5682    150k   Same
    5683     29k   Same
    
    This completed the MRS goals at 90deg, so I moved it to 
    40deg for a 0-field run:
    5684     45k   MRS@40, D5 OFF. FS as before.
    
    After finishing this I turned MRS to 45deg and switched
    D5 to A polarity.
    5685     49k   MRS@45, D5@4A. FS as before.
            ----
            399k trigger 6 :-)
    
    5686 Pedestal run
    5687 TPC pedestal run
    5688 DriftVelMon pedestal run
    
    3: I've seen no detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Nov 06 2001 - 10:23:15 EST